# Mixed Planewave and Localized Orbital Basis for Sparse-Stochastic Hybrid TDDFT

### v1

@@ Acknowledgements.
Figure 8: LUMO-HOMO LDA gap differences (compared to the PW-LDA result) in eV calculated for C20H10, using various AO basis sets. The numbers in parentheses indicate the total number of basis functions used. The minimum deviation AO set is aug-cc-pVDZ with 550 total basis functions, where the gap difference between AO and PW is $-6.05\times 10^{-4}$ eV.
Localized AO basis sets vary widely in their composition and completeness, ranging from minimal sets like STO-3G to highly extensive ones such as aug-cc-pVQZ. In principle, as an AO basis set becomes more complete, its results should asymptotically approach those of PW calculations. However, this increased completeness comes at a significantly higher computational cost. To identify an optimal AO basis set for the mixed-basis scheme, we benchmark several AO basis sets by performing LDA-DFT calculations and comparing the resulting LUMO-HOMO gaps against those from a reference PW LDA-DFT calculation. The error is defined as the difference between AO and the reference PW gaps. All benchmarks are conducted on the C20H10 (Fig. 8).
Since our calculations employ real-space grids, we require basis functions with smooth spatial decay to maintain numerical stability in integrations.Becke (1988a, b) The aug-cc-pVDZ basis set satisfies this need by avoiding excessively sharp Gaussian exponents ($\zeta$): carbon’s tightest exponent in aug-cc-pVDZ has $\zeta\sim 8\times 10^{3}$ Bohr-2 is far smoother than in sharper sets like 6-31G with $\zeta\sim 2\times 10^{4}$ Bohr-2 or STO-3G with $\zeta\sim 7\times 10^{4}$ Bohr-2.Dunning (1989); Pritchard et al. (2019) This careful balance prevents the grid artifacts that would arise from extreme exponents while maintaining accuracy for both core and valence electrons.
@@
Appendix C: Computational Parameters
@@
System $N_{x}$ $N_{y}$ $N_{z}$ $dx$ $N_{v}$ $N_{c}$ $N_{k_{\text{low}}}$
C20H10 70 70 40 0.4 45 $45-160$ 201
Chla 106 96 80 0.4 116 $174-464$ 867
ICG-7 130 100 80 0.4 115 $173-460$ 1127
Flav-9 156 100 70 0.4 116 $174-464$ 1195
Table 4: Computational parameters for the four test systems: grid extents, isotropic grid-spacing of $dx=dy=dz=0.4$ Bohr is used for all systems, $N_{v}$ values, and $N_{c}$ ranges used for sparse-stochastic hybrid DFT and TDDFT calculations. $N_{k_{\text{low}}}$ is the number of deterministically treated long-wavelength terms in the exchange kernel $u(k)$. The high-$k$ space is represented with 1000 sparse-stochastic vectors, details are provided in Refs. Bradbury et al. (2023a); Sereda et al. (2024).
Data Availability
The data that supports the findings of this study are available within the article and appendixes. Additional data that support the findings of this study are available from the corresponding author upon request.

### v2

@@ 4 Conclusion
Figure 8: LUMO-HOMO LDA gap differences (compared to the PW-LDA result) in eV calculated for C20H10, using various AO basis sets. The numbers in parentheses indicate the total number of basis functions used. The minimum deviation AO set is aug-$c$-pVDZ with 550 total basis functions, where the gap difference between AO and PW is $6.05\times 10^{-4}$ eV.
Localized AO basis sets vary widely in their composition and completeness, ranging from minimal sets like STO-3G to highly extensive ones such as aug-cc-pVQZ. In principle, as an AO basis set becomes more complete, its results should asymptotically approach those of PW calculations. However, this increased completeness comes at a significantly higher computational cost. To identify an optimal AO basis set for the mixed-basis scheme, we benchmark several AO basis sets by performing LDA-DFT calculations and comparing the resulting LUMO-HOMO gaps against those from a reference PW LDA-DFT calculation. The error is defined as the difference between AO and the reference PW gaps. All benchmarks are conducted on the C20H10 (Fig. 8). Table 3 demonstrates that static gaps (calculated via our PW/AO sparse-stochastic hybrid-DFT with CAM-LDA0 functional) of C20H10 exhibit minimal sensitivity ($\leq 0.037$ eV variation) across nine different AO basis sets. The aug-cc-pVDZ basis performs particularly well, differing by only 0.009 eV from the PW reference.
Basis Set Static Gap (eV) Abs. Difference (eV)
Pure PW 6.712 (Reference)
aug-$c$-pVDZ 6.704 0.009
aug-$c$-pVTZ 6.703 0.009
aug-$c$-pVQZ 6.701 0.012
3-21G* 6.694 0.018
$c$-pVDZ 6.680 0.032
def2-SVP 6.678 0.034
STO-3G 6.677 0.035
6-311G* 6.675 0.037
Table 3: Static LUMO-HOMO gap (in eV) of C20H10 calculated via PW/AO sparse-stochastic hybrid-DFT with CAM-LDA0 functional for various AO basis sets compared to the pure PW reference.
Since our calculations employ real-space grids, we require basis functions with smooth spatial decay to maintain numerical stability in integrations.55, 56 The aug-cc-pVDZ basis set satisfies this need by avoiding excessively sharp Gaussian exponents ($\zeta$): carbon’s tightest exponent in aug-cc-pVDZ has $\zeta\sim 8\times 10^{3}$ Bohr-2 is far smoother than in sharper sets like 6-31G with $\zeta\sim 2\times 10^{4}$ Bohr-2 or STO-3G with $\zeta\sim 7\times 10^{4}$ Bohr-2.57, 58 This careful balance prevents the grid artifacts that would arise from extreme exponents while maintaining accuracy for both core and valence electrons.
@@
Appendix C: Real-space Grid and Computational Parameters
@@
System $N_{x}$ $N_{y}$ $N_{z}$ dxdx $N_{v}$ $N_{c}$ $N_{k_{\text{low}}}$
C20H10 70 70 40 0.4 45 4516045-160 201
Chla 106 96 80 0.4 116 174464174-464 867
ICG-7 130 100 80 0.4 115 173460173-460 1127
Flav-9 156 100 70 0.4 116 174464174-464 1195
Table 5: Computational parameters for the four test systems: grid extents, isotropic grid-spacing of $dx=dy=dz=0.4$ Bohr is used for all systems, $N_{v}$ values, and $N_{c}$ ranges used for sparse-stochastic hybrid DFT and TDDFT calculations. $N_{k_{\text{low}}}$ is the number of deterministically treated long-wavelength terms in the exchange kernel $u(k)$. The high-$k$ space is represented with 1000 sparse-stochastic vectors, details are provided in Refs. 24, 25.
Fig. 9 illustrates an example for the real-space grid (“box”) configuration for the C20H10 molecule, employing a uniform grid spacing of 0.4 Bohr. The molecule is centered at the origin, and the grid extends along each Cartesian direction with a minimum 6 Bohr padding to ensure negligible boundary effects. Identical criteria (fixed spacing + padding) are applied to other molecules, with box sizes adjusted to accommodate their respective dimensions.
Figure 9: C20H10 molecule (black balls) and the overall simulation grid (gray dots) viewed from different directions.